OpenStudy (anonymous):

divide the polynomials. ( -12x^8 + x^6 - 8x^3) divided by (4x^2)

OpenStudy (anonymous):

use synthetic division?

OpenStudy (anonymous):

Factor the numerator and then simplify.\[\frac{-12 x^8+x^6-8 x^3}{4 x^2}=\frac{-x^3 \left(12 x^5-x^3+8\right)}{4 x^2}=-\frac{1}{2} x \left(12 x^5-x^3+8\right) \]

OpenStudy (anonymous):

It's -1/4 x (12x^5 - x^3 + 8) ! Check his or her work, even though it is by a high level mathematician!

OpenStudy (anonymous):

thecookiemonster is correct. The 1/2 should be changed to 1/4.